# Refining Concentration for Gaussian Quadratic Chaos

### v1

@@ III Proof of Proposition 1
where $\boldsymbol{\xi}_{1},\cdots,\boldsymbol{\xi}_{n}$ are independent standard normal random variables that are the entries of the random vector $\underline{\boldsymbol{\xi}}=U^{T}\underline{\boldsymbol{x}}$. Then
@@
Using $\mathbb{E}[e^{s\lambda_{i}\boldsymbol{\xi}^{2}_{i}}]=\frac{1}{\sqrt{1-2s\lambda_{i}}}$ for $12s\lambda_{i}<$ 1, we get
@@
for all $s\geq 0$ such that $12s\lambda_{i}<$ 1 for all $1\leq i\leq n$. Now, let $a,b>0$ be such that (6) holds. Then the bound in (32) can be loosened as
@@
for all $s\geq 0$ such that $|2s\lambda_{i}|\leq b$ for all $1\leq i\leq n$, or equivalently, $0\leq s\leq\frac{b}{2\|A\|}$. The minimum for the function
@@
over $0\leq s\leq\frac{b}{2\|A\|}$ occurs at $s_{opt}=\min\{s_{0},\frac{b}{2\|A\|}\}$ where $s_{0}=\frac{t}{4a\|A\|_{2}^{2}}$. We consider two cases:
@@
Next, let us explore the constants $a,b>0$. Reference [3] provides $a=1$ and $b=\frac{1}{2}$. This results in $\kappa=\kappa(1,0.5)=0.125$. But, one can do better. We need the special case of the following lemma for $m=1$:
@@
Let $m\geq 1$ be an integer. For every $a>\frac{1}{m+1}$, the largest $b$ for which the inequality in (6) holds is $\theta_{m}^{-1}(a)$ where $\theta_{m}^{-1}(\cdot)$ is the inverse function to $\theta_{m}(\cdot)$ defined in (7). Equivalently, for every $10<b<$ 1, the smallest $a$ (tightest upper bound) for which (6) holds is $a=\theta_{m}(b)$. Conversely, if $a\leq\frac{1}{m+1}$, there exists no $b>0$ such that (6) holds.
@@
Using Lemma 1 with $m=1$ results in
@@
for arbitrary $10<b<$ 1. Since $\frac{1}{8\theta_{1}(b)}$ is decreasing in $b$ and $\frac{b}{4}$ is increasing in $b$, the largest value for their minimum is achieved when $\frac{1}{8\theta_{1}(b)}=\frac{b}{4}$, i.e., $12b\theta_{1}(b)=$ 1. This is solved for $b\approx 0.583$ and the achieved maximum value is approximately $0.1457$ as shown in Fig. 1
Figure 1: The maximum for $\kappa(\theta_{1}(b),b)=\min\Big{\{}\frac{1}{8\theta_{1}(b)},\frac{b}{4}\Big{\}}$ is achieved where $\frac{1}{8\theta_{1}(b)}=\frac{b}{4}$. This is solved for $b\approx 0.583$ and the achieved maximum value is approximately $0.1457$.
@@ IV Proof of Proposition 2
for all $s\geq 0$ such that $0\leq 2s\lambda_{i}\leq b$, or equivalently, $0\leq s\leq\frac{b}{2\|A\|}$. Since $ab<1$ by assumption, then
@@
At this point, we adopt the notations
$\alpha=\|A\|,\,\,\,\,\,\beta=\|A\|_{2}^{2},$ (42)
for notational simplicity. Let us call the exponent on the right side of (41) by $f(s)$, i.e.,
@@
It is straightforward to check that the function $f$ is convex and achieves its minimum over its domain $[0,\frac{1}{2a\alpha})$ at
@@
It follows that the expression on the right side of (41) achieves its minimum at
@@
    $f(s_{opt})=f\Big{(}\frac{b}{2\|A\|}\Big{)}=-\frac{bt}{2\alpha}+\frac{b^{2}\beta}{4(1-ab)\alpha^{2}}$ (48)
@@
To recap, we have shown that
@@
Next, the next lemma explores a possibility for $a,b$. It addresses an inequality for which (10) is a special case.
@@
Using Lemma 2 for $m=1$, we obtain
@@ V Proof of Proposition 3
$\Pr(\boldsymbol{\Delta}>t)\leq e^{-\kappa\min\left\{\frac{t^{2}}{\beta},\frac{t}{\alpha}\right\}}.$ (58)
Note
@@
We aim to show that there is $\kappa>0$ such that the upper bound in (50) is less than or equal to the upper bound in (58), or equivalently, $f(s_{opt})\leq-\kappa\min\{\frac{\beta}{t^{2}},\frac{\alpha}{t}\}$. We want this to hold regardless of $t,\alpha,\beta$. Thus, we look for largest $\kappa>0$ such that
$\kappa\leq\inf_{t,\alpha,\beta>0}-\frac{f(s_{opt})}{\min\{\frac{t^{2}}{\beta},\frac{t}{\alpha}\}}.$ (62)
Interestingly, the ratio $-\frac{f(s_{opt})}{\min\{\frac{t^{2}}{\beta},\frac{t}{\alpha}\}}$ depends on $t,\alpha,\beta$ only through the ratio
$\rho=\frac{\alpha t}{\beta}.$ (63)
For ease of comparing the right sides in (16) and (61), we fix $a$ such that $c=1$, i.e., $\frac{3a-2}{2(1-a)^{2}}=1$. This quadratic equation has only one root in the admissible domain $\frac{2}{3}<a<1$. We denote this value of $a$ by $a_{0}$ given by
@@
One can easily check that the corresponding value for $b$ by $\frac{3a_{0}-2}{a_{0}(2a_{0}-1)}=\frac{1}{2}$. Then we have
$-\frac{f(s_{opt})}{\min\{\frac{t^{2}}{\beta},\frac{t}{\alpha}\}}=\left\{\begin{array}[]{cc}\frac{1}{2a_{0}\rho}-\frac{1}{2a_{0}^{2}\rho^{2}}\big{(}\sqrt{1+2a_{0}\rho}-1\big{)}&\rho\leq 1\\ \frac{1}{4}-\frac{1}{16(1-\frac{a_{0}}{2})\rho}&\rho>1\end{array}\right..$ (67)
Fig. 2 shows the graph of the function of $\rho$ on the right side of (67). It achieves its absolute minimum value at $\rho=1$. It follows that the value for the infimum on the right side of (62) is
@@
Figure 2: Graph of the function of $\rho$ on the right side of (67). It achieves its absolute minimum value at $\rho=1$.
It is shown in Appendix C that the choice of $a=a_{0}$ in (64) results in the largest value for $\kappa$.
Finally, we address LM inequality itself. Let us write LM inequality in (4) slightly differently as follows. Call $2\sqrt{\beta t}+2\alpha t=t^{\prime}$ where $\alpha,\beta$ are as in (42). Solving for $t$, we get $t=(\frac{\sqrt{\beta+2\alpha t^{\prime}}-\sqrt{\beta}}{2\alpha})^{2}$. Relabeling $t^{\prime}$ as $t$, LM inequality becomes
@@
Consider HW inequality in (58). LM inequality implies HW inequality for all $\kappa$ satisfying
$\kappa\leq\inf_{t,\alpha,\beta>0}-\frac{(\frac{\sqrt{\beta+2\alpha t}-\sqrt{\beta}}{2\alpha})^{2}}{\min\{\frac{t^{2}}{\beta},\frac{t}{\alpha}\}}.$ (70)
Once again, we realize that the ratio on the right side of (70) depends entirely on the ratio $\rho=\frac{\alpha t}{\beta}$. It is given by
@@
One can easily check that this function of $\rho$ achieves it absolute minimum value of $1-\frac{\sqrt{3}}{2}$ at $\rho=1$. This establishes Proposition 3.
@@ VI Proof of Proposition 4
$=$ $\exp\Big{(}-st+a2^{m}s^{m+1}\sum_{i=1}^{n}|\lambda_{i}|^{m+1}+\sum_{k=2}^{m}\frac{1}{k}2^{k-1}s^{k}\sum_{i=1}^{n}\lambda_{i}^{k}\Big{)}$
@@
for all $s\geq 0$ such that $2s\lambda_{i}\leq b$, or equivalently, $0\leq s\leq\frac{b}{2\|A\|}$. The value of $s$ that minimizes the right side of (74) in terms of $s$ is not tractable for arbitrary integer $m$. Instead, let $s_{0}$ be the value of $s$ such that the sum of the first two terms in the exponent, i.e., $-st+a2^{m}s^{m+1}\|A\|_{m+1}^{m+1}$ is minimized. It is given by
$s_{0}=\frac{1}{2}\left(\frac{t}{a(m+1)\|A\|_{m+1}^{m+1}}\right)^{\frac{1}{m}}.$ (75)
We plug in $s=s_{*}$ on the right side in (74) where333
$s_{*}=\min\Big{\{}s_{0},\frac{b}{2\|A\|}\Big{\}}.$ (76)
The third term in the exponent in (74) is bounded by
$\sum_{k=2}^{m}\frac{2^{k-1}s_{*}^{k}\mathrm{Tr}(A^{k})}{k}$$\leq$ $\sum_{k=2}^{m}\frac{2^{k-1}(\frac{b}{2\|A\|})^{k}\mathrm{Tr}(A^{k})}{k}$ (77)
$=$ $\sum_{k=2}^{m}\frac{b^{k}}{2k}\frac{\mathrm{Tr}(A^{k})}{\|A\|^{k}}$
$\leq$ $\frac{n}{2}\sum_{k=2}^{m}\frac{b^{k}}{k},$
where the last step is due to $\mathrm{Tr}(A^{k})\leq n\|A^{k}\|\leq n\|A\|^{k}$. Next, we bound the sum of the first two terms, namely, $-s_{*}t+a2^{m}s_{*}^{m+1}\|A\|_{m+1}^{m+1}$. We consider two cases:
@@
    Let $s_{0}\leq\frac{b}{2\|A\|}$. Then $s_{*}=s_{0}$ and simple algebra shows that
    $-s_{*}t+a2^{m}s_{*}^{m+1}\|A\|_{m+1}^{m+1}=-\kappa_{m,1}\frac{t^{1+\frac{1}{m}}}{\|A\|_{m+1}^{1+\frac{1}{m}}}.$ (78)
    where
    $\kappa_{m,1}=\frac{1-\frac{1}{m+1}}{2a^{\frac{1}{m}}(m+1)^{\frac{1}{m}}}$ (79)
@@
    Let $s_{0}>\frac{b}{2\|A\|}$. Then $s_{opt}=\frac{b}{2\|A\|}$ and we have
    $-s_{*}t+a2^{m}s_{*}^{m+1}\|A\|_{m+1}^{m+1}$ $=$ $-s_{*}t+a2^{m}s_{*}\times s_{*}^{m}\|A\|_{m+1}^{m+1}$ (80)
@@
    $=$ $-\frac{bt}{2\|A\|}+a2^{m}\frac{b}{2\|A\|}\times\frac{t}{a2^{m}(m+1)\|A\|_{m+1}^{m+1}}\|A\|_{m+1}^{m+1}$
@@
    where
@@
$-s_{opt}t+a2^{m}s_{opt}^{m+1}\mathrm{tr}(A^{m+1})$$\leq$ $\max\Big{\{}-\kappa_{m,1}\frac{t^{1+\frac{1}{m}}}{\|A\|_{m+1}^{1+\frac{1}{m}}},-\kappa_{m,2}\frac{t}{\|A\|}\Big{\}}$ (82)
$=$ $-\min\Big{\{}\kappa_{m,1}\frac{t^{1+\frac{1}{m}}}{\|A\|_{m+1}^{1+\frac{1}{m}}},\kappa_{m,2}(a,b)\frac{t}{\|A\|}\Big{\}}$

### v2

@@ III Proof of Proposition 1
where $\boldsymbol{\xi}_{1},\cdots,\boldsymbol{\xi}_{n}$ are independent standard normal random variables that are the entries of the random vector $\underline{\boldsymbol{\xi}}=U^{T}\underline{\boldsymbol{x}}$. By independence,
@@
Since $\mathbb{E}[e^{s\lambda_{i}\boldsymbol{\xi}^{2}_{i}}]=\frac{1}{\sqrt{1-2s\lambda_{i}}}$ for $12s\lambda_{i}<$ 1, we get
@@
for every $s\geq 0$ such that $12s\lambda_{i}<$ 1 for all $1\leq i\leq n$. Let $a,b>0$ be such that (7) holds. The bound in (99) can be loosened as
@@
for every $s\geq 0$ such that $|2s\lambda_{i}|\leq b$ for all $1\leq i\leq n$, or equivalently, $0\leq s\leq\frac{b}{2\|A\|}$. The minimum value for the function
@@
over $0\leq s\leq\frac{b}{2\|A\|}$ occurs at $s_{opt}=\min\{s_{0},\frac{b}{2\|A\|}\}$ where $s_{0}=\frac{t}{4a\|A\|_{2}^{2}}$. We consider two cases.
@@
Next, let us explore the constants $a,b>0$. Reference [3] considers $a=1$ and $b=\frac{1}{2}$. This results in $\kappa=\kappa(1,0.5)=0.125$. But, one can do better. We need the special case of the following lemma for $m=1$.
@@
Let $m\geq 1$ be an integer. For every $10<b<$ 1, the smallest $a$ (tightest upper bound) for which (7) holds is $a=\theta_{m}(b)$ as defined in (8). Equivalently, for every $a>\frac{1}{m+1}$, the largest $b$ for which the inequality in (7) holds is $\theta_{m}^{-1}(a)$ where $\theta_{m}^{-1}(\cdot)$ is the inverse function to $\theta_{m}(\cdot)$. Conversely, if $a\leq\frac{1}{m+1}$, there exists no $b>0$ such that (7) holds.
@@
Applying Lemma 1 with $m=1$ results in
@@
for arbitrary $10<b<$ 1. Since $\frac{1}{8\theta_{1}(b)}$ is decreasing in $b$ and $\frac{b}{4}$ is increasing in $b$, the largest value for their minimum is achieved when $\frac{1}{8\theta_{1}(b)}=\frac{b}{4}$, i.e., $12b\theta_{1}(b)=$ 1. This equation is solved for $b\approx 0.583$ and the achieved maximum value is approximately $0.1457$ as shown in Fig. 10.
Figure 10: The maximum value for $\kappa(\theta_{1}(b),b)=\min\big{\{}\frac{1}{8\theta_{1}(b)},\frac{b}{4}\big{\}}$ is achieved where $\frac{1}{8\theta_{1}(b)}=\frac{b}{4}$. This is solved for $b\approx 0.583$ and the attained maximum value is approximately $0.1457$.
@@ IV Proof of Proposition 2
for every $s\geq 0$ such that $0\leq 2s\lambda_{i}\leq b$ for all $1\leq i\leq n$, or equivalently, $0\leq s\leq\frac{b}{2\|A\|}$. Since $ab<1$ by assumption, then
@@
Let $\alpha$ and $\beta$ be as in (12) and call the exponent on the right side of (108) by $f(s)$, i.e.,
@@
It is straightforward to check that the function $f$ is convex (concave upward) over the entire interval $(-\infty,\frac{1}{2a\alpha})$ on the left side of its vertical asymptote $s=\frac{1}{2a\alpha}$. Moreover, it achieves an absolute minimum value over that interval at
@@
Note that $\frac{b}{2\alpha}<\frac{1}{2a\alpha}$ due to the assumption $ab<1$. It follows that the expression on the right side of (108) achieves its minimum value over $0\leq s\leq\frac{b}{2\alpha}$ at
@@
    $f(s_{opt})=f\Big{(}\frac{b}{2\alpha}\Big{)}=-\frac{bt}{2\alpha}+\frac{b^{2}\beta}{4(1-ab)\alpha^{2}}$ (114)
@@
We have shown that
@@
The next lemma explores possibilities for $a,b$. It addresses an inequality for which (11) is a special case.
@@
Applying Lemma 2 with $m=1$, we obtain
@@ V Proof of Proposition 3
$\Pr(\boldsymbol{\Delta}>t)\leq e^{-\kappa\min\left\{\frac{t^{2}}{\beta},\frac{t}{\alpha}\right\}},$ (124)
where $\alpha,\beta$ are given in (12). Note
@@
We aim to show that there exists $\kappa>0$ such that the upper bound in (14) is less than or equal to the upper bound in (124), or equivalently, $\Lambda(t,a)\leq-\kappa\min\{\frac{t^{2}}{\beta},\frac{t}{\alpha}\}$. We want this to hold regardless of $t,\alpha,\beta$ and $\kappa$ to be as large as possible. Thus, we are looking to compute
$\kappa=\inf_{t,\alpha,\beta>0}-\frac{\Lambda(t,a)}{\min\{\frac{t^{2}}{\beta},\frac{t}{\alpha}\}}.$ (128)
Interestingly, $-\frac{\Lambda(t,a)}{\min\{\frac{t^{2}}{\beta},\frac{t}{\alpha}\}}$ depends on $t,\alpha,\beta$ only through the ratio $\rho=\frac{\alpha t}{\beta}$ in (36). For convenience in comparing the right sides in (17) and (127), we fix $a$ such that $c=1$, i.e., $\frac{3a-2}{2(1-a)^{2}}=1$. This quadratic equation has only one root in the admissible interval $\frac{2}{3}<a<1$. We denote this value of $a$ by $a_{0}$ given by
@@
One can easily check that the corresponding value for $b$ is $b_{0}=\frac{3a_{0}-2}{a_{0}(2a_{0}-1)}=\frac{1}{2}$. We have
$-\frac{\Lambda(t,a_{0})}{\min\{\frac{t^{2}}{\beta},\frac{t}{\alpha}\}}=\left\{\begin{array}[]{cc}\frac{1}{2a_{0}\rho}-\frac{1}{2a_{0}^{2}\rho^{2}}\big{(}\sqrt{1+2a_{0}\rho}-1\big{)}&\rho\leq 1\\ \frac{1}{4}-\frac{1}{16(1-\frac{a_{0}}{2})\rho}&\rho>1\end{array}\right..$ (132)
Fig. 11 shows the graph of the function of $\rho$ that sits on the right side of (132). It achieves its absolute minimum value at $\rho=1$. It follows that the infimum in (128) is given by
@@
Figure 11: Graph of the function of $\rho$ given on the right side of (132). It achieves its absolute minimum value at $\rho=1$.
It is shown in Appendix C that the choice of $a=a_{0}$ in (129) indeed results in the largest value for $\kappa$ and that it was not just a choice of convenience.
Finally, we address LM inequality itself. To see why (4) and (5) are equivalent, call $t^{\prime}=2\sqrt{\beta t}+2\alpha t$. Solving for $t$, we get $t=(\frac{\sqrt{\beta+2\alpha t^{\prime}}-\sqrt{\beta}}{2\alpha})^{2}$. Relabeling $t^{\prime}$ as $t$, LM inequality in (4) becomes
@@
This inequality implies HW inequality in (124) with
$\kappa=\inf_{t,\alpha,\beta>0}-\frac{(\frac{\sqrt{\beta+2\alpha t}-\sqrt{\beta}}{2\alpha})^{2}}{\min\{\frac{t^{2}}{\beta},\frac{t}{\alpha}\}}.$ (135)
Once again, the ratio on the right side of (135) depends entirely on $\rho=\frac{\alpha t}{\beta}$. It is given by
@@
This function of $\rho$ achieves it absolute minimum value of $1-\frac{\sqrt{3}}{2}$ at $\rho=1$.
@@ VI Proof of Proposition 4
$=$ $\exp\Big{(}-st+a2^{m}s^{m+1}\sum_{i=1}^{n}|\lambda_{i}|^{m+1}+\sum_{k=2}^{m}\frac{2^{k-1}s^{k}}{k}\sum_{i=1}^{n}\lambda_{i}^{k}\Big{)}$
@@
for every $s\geq 0$ such that $|2s\lambda_{i}|\leq b$ for all $1\leq i\leq n$, or equivalently, $0\leq s\leq\frac{b}{2\|A\|}$. The value of $s$ that minimizes the right side of (139) is not tractable for arbitrary integer $m$. Instead, let $s_{0}$ be the value of $s$ such that the sum of the first two terms in the exponent, i.e., $-st+a2^{m}s^{m+1}\|A\|_{m+1}^{m+1}$ is minimized. It is given by
$s_{0}=\frac{1}{2}\left(\frac{t}{(m+1)a\|A\|_{m+1}^{m+1}}\right)^{\frac{1}{m}}.$ (140)
We compute the right side of (139) at $s=\hat{s}_{opt}$ defined by
$\hat{s}_{opt}=\min\Big{\{}s_{0},\frac{b}{2\|A\|}\Big{\}}.$ (141)
Then the third term in the exponent in (139) is bounded by
$\sum_{k=2}^{m}\frac{2^{k-1}\hat{s}_{opt}^{k}\mathrm{Tr}(A^{k})}{k}\leq\sum_{k=2}^{m}\frac{2^{k-1}(\frac{b}{2\|A\|})^{k}|\mathrm{Tr}(A^{k})|}{k}=\sum_{k=2}^{m}\frac{b^{k}}{2k}\frac{|\mathrm{Tr}(A^{k})|}{\|A\|^{k}}\leq\frac{n}{2}\sum_{k=2}^{m}\frac{b^{k}}{k},$ (142)
where the first step is due to $\hat{s}_{opt}^{k}\mathrm{Tr}(A^{k})\leq|\hat{s}_{opt}^{k}\mathrm{Tr}(A^{k})|$ and $|\hat{s}_{opt}|=\hat{s}_{opt}\leq\frac{b}{2\|A\|}$ and the last step is due to [^17] For a symmetric matrix $A$, the operator norm $\|A\|$ is the maximum value among the absolute values of eigenvalues of $A$.
$|\mathrm{Tr}(A^{k})|=\big{|}\sum_{i=1}^{n}\lambda_{i}^{k}\big{|}\leq\sum_{i=1}^{n}|\lambda_{i}|^{k}\leq n(\max_{1\leq i\leq n}|\lambda_{i}|)^{k}=n\|A\|^{k}.$ (143)
Next, we bound the sum of the first two terms in the exponent, i.e., $-\hat{s}_{opt}t+a2^{m}\hat{s}_{opt}^{m+1}\|A\|_{m+1}^{m+1}$. We consider two cases.
@@
    Let $s_{0}\leq\frac{b}{2\|A\|}$. Then $\hat{s}_{opt}=s_{0}$ and simple algebra shows that
    $-\hat{s}_{opt}t+a2^{m}\hat{s}_{opt}^{m+1}\|A\|_{m+1}^{m+1}=-\kappa_{m,1}\frac{t^{1+\frac{1}{m}}}{\|A\|_{m+1}^{1+\frac{1}{m}}},$ (144)
    where $\kappa_{m,1}$ is defined by
    $\kappa_{m,1}=\frac{1-\frac{1}{m+1}}{2(m+1)^{\frac{1}{m}}a^{\frac{1}{m}}}.$ (145)
@@
    Let $s_{0}>\frac{b}{2\|A\|}$. Then $\hat{s}_{opt}=\frac{b}{2\|A\|}$ and we have
    $-\hat{s}_{opt}t+a2^{m}\hat{s}_{opt}^{m+1}\|A\|_{m+1}^{m+1}$ $=$ $-\hat{s}_{opt}t+a2^{m}\hat{s}_{opt}\times\hat{s}_{opt}^{m}\|A\|_{m+1}^{m+1}$ (146)
@@
    $=$ $-\frac{bt}{2\|A\|}+a2^{m}\frac{b}{2\|A\|}\times\frac{t}{2^{m}(m+1)a\|A\|_{m+1}^{m+1}}\|A\|_{m+1}^{m+1}$
@@
    where the second step is due to $\hat{s}_{opt}\leq s_{0}$ and $\kappa_{m,2}$ is defined by
@@
$-\hat{s}_{opt}t+a2^{m}\hat{s}_{opt}^{m+1}\|A\|_{m+1}^{m+1}$$\leq$ $\max\Big{\{}-\kappa_{m,1}\frac{t^{1+\frac{1}{m}}}{\|A\|_{m+1}^{1+\frac{1}{m}}},-\kappa_{m,2}\frac{t}{\|A\|}\Big{\}}$ (148)
$=$ $-\min\Big{\{}\kappa_{m,1}\frac{t^{1+\frac{1}{m}}}{\|A\|_{m+1}^{1+\frac{1}{m}}},\kappa_{m,2}\frac{t}{\|A\|}\Big{\}}$